OpenStudy (anonymous):

James has 33 coins in his pocket, all of them nickels and quarters. If he has a total of 3.25, how many quarters does he have?

OpenStudy (brooke2353):

Let's say that James has "n" nickels and "q" quarters. From the given number of coins: n + q = 33 .......... (1) From the coin values and the total amount (in cents): 5n + 25q = 325 ....... (2) Using the elimination method ... (2) – 5(1): 5n + 25q = 325 – 5n + 5q = 165 –––––––––––– 20q = 160 q = 8 James has 8 quarters in his pocket. (He would also have 25 nickels. Total values: $2.00 + $1.25 = $3.25)

OpenStudy (anonymous):

wow! Thank you very much!

OpenStudy (brooke2353):

your welcome